Use the starting balance sheet and the list of changes to create an updated balance sheet and to answer the question.

Hopewell Corporation
Balance Sheet
As of December 31, 2018
(amounts in thousands)
Cash 2,900 Liabilities 2,400
Other Assets 3,700 Equity 4,200
Total Assets 6,600 Total Liabilities & Equity 6,600

Between January 1 and March 31, 2019:

1. Cash decreases by $300,000
2. Liabilities increase by $400,000
3. Equity decreases by $200,000

What is the value for Other Assets on March 31, 2019?

Please specify your answer in the same units as the balance sheet.
Answer #1

Solution:

Inregards to the information mentioned in the question following the value of other assets on March 31,2019

Effect of transactions between Jan 1 and March 31,2019

1. Cash decreases by $300,000-

Cash = 2900-300=2600
2. Liabilities increase by $400,000

Liabilities = 2400+400=2800
3. Equity decreases by $200,000

Equity = 4200-200=4000

Hopewell Corporation
Balance Sheet
As of March 31, 2019
(amounts in thousands)
Cash 2600 Liabilities 2,800
Other Assets 4200 Equity 4000
Total Assets 6,800 Total Liabilities & Equity 6 800

As s result of all the above transaction:

Value of Other Assets on March 31,2019

=Total Assets - Cash

= 6800-2600

=4200